use the given matrix to perform the operation R1+2R2 and select the correct option.
3, -2, 1
1, -3, 4
1, -2, 0
a. 3, -2, 1
5, -8, 9
1, -2, 0
b. 3, -2, 1
2, -6, 8
13, -2, 0
c. 3, -2, 0
1, -3, 4
1/5, -2, 0
d. 5, -8, 9
1, -3, 4
1, -2, 0

Answer by stanbon(75887)

use the given matrix to perform the operation R1+2R2 and select the correct option.
3, -2, 1
1, -3, 4
1, -2, 0
------------
3+2*1 = 5 ; -2+2*-3 = -8 ; 1+2*4 = 9
-------------------
The answer is,
d. 5, -8, 9
1, -3, 4
1, -2, 0
--------------
Cheers,
Stan H.
-------------